OpenStudy (anonymous):

Eliminate the “z” variable in the first two equations by finding opposite coefficients and adding. What is the resulting equation? 4x – y + z = 6 3x + 2y – z = 0 4x – y – 3z = 2

OpenStudy (amistre64):

they already HAVE opposite coeffs

OpenStudy (amistre64):

so just add them together :/

OpenStudy (anonymous):

x + y = 6 7x + y = 6 7x – y = 6 x – y = 6 these are the answers ...

OpenStudy (amistre64):

those are the options, not the answers; only one of the options is the answer .....

OpenStudy (amistre64):

add the first 2 eqs and what do we get?

OpenStudy (anonymous):

7x+y=6

OpenStudy (amistre64):

4x – y + z = 6 3x + 2y – z = 0 -------------- 7x +1y -0z = 6

OpenStudy (amistre64):

I agree :)